OpenStudy (anonymous):

What is f^7 (2) in this series?

OpenStudy (anonymous):

f(x) = \[\sum_{n=0}^{\infty} (-1)^{n} \sqrt{n} (x-2)^{n} / n!\]

OpenStudy (irishboy123):

what is x - 2 if x = 2?

OpenStudy (anonymous):

0... Yes I put in 0 but it didn't work...

OpenStudy (thomas5267):

Is f^7(x) the the seventh derivative (i.e. \(f^{(7)}(x)\)) of this series or \((f(x))^7\)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

it is the seventh derivative

OpenStudy (thomas5267):

It is not zero because the 8th term of f(x) will be a constant when differentiated 7th time.

OpenStudy (thomas5267):

How do I know that? Mathematica!

OpenStudy (anonymous):

Wait could you show me that?

OpenStudy (perl):

is this the series

OpenStudy (anonymous):

yes

OpenStudy (perl):

$$ \Large f(x) = \sum_{n=0}^{\infty} \frac{(-1)^{n} \sqrt{n}~ (x-2)^{n}}{ n!} $$

OpenStudy (anonymous):

yup!!

OpenStudy (perl):

$$ \Large { f(x) = \sum_{n=0}^{\infty} \frac{(-1)^{n} ~\sqrt{n}~ (x-2)^{n}}{ n!} \\~\\ f ' (x) = \sum_{n=1}^{\infty} \frac{(-1)^{n} \sqrt{n}~\cdot n\cdot (x-2)^{n-1}}{ n!} } $$

OpenStudy (perl):

lets use rational exponents

OpenStudy (perl):

do you see the pattern emerging>?

OpenStudy (thomas5267):

\[ \begin{align*} &\quad\sum_{n=0}^9 \frac{(-1)^{n}\sqrt{n}(x-2)^{n}}{ n!}\\ &=-(x-2)+\frac{(x-2)^2}{\sqrt{2}}-\frac{(x-2)^3}{2 \sqrt{3}}+\frac{1}{12}(x-2)^4\\ &\quad -\frac{(x-2)^5}{24 \sqrt{5}}+\frac{(x-2)^6}{120 \sqrt{6}}-\frac{(x-2)^7}{720 \sqrt{7}}+\frac{(x-2)^8}{10080 \sqrt{2}}-\frac{(x-2)^9}{120960} \end{align*} \] After differentiating it 7 times, the first 7 terms vanishes, the 8th term will become a constant and the 9+ term will have (x-2) as a term. Hence, you only have to differentiate the 8th term 7 times to get \(f^{(7)}(2)\).

OpenStudy (thomas5267):

Wait no! The 7th term will become a constant and the 8+ terms will have (x-2) as a term.

OpenStudy (perl):

do you see why it canceled?

OpenStudy (perl):

$$ \Large { f(x) = \sum_{n=0}^{\infty} \frac{(-1)^{n} ~{n}^{1/ 2}~ (x-2)^{n}}{ n!} \\~\\\text{power rule}\\ f ' (x) = \sum_{n=1}^{\infty} \frac{(-1)^{n} ~{n}^{1/ 2}~\cdot n\cdot (x-2)^{n-1}}{ n!} \\~\\=\sum_{n=1}^{\infty} \frac{(-1)^{n} ~{n}^{1/ 2}~\cdot n\cdot (x-2)^{n-1}}{ n\cdot (n-1)!} \\~\\ =\sum_{n=1}^{\infty} \frac{(-1)^{n} ~{n}^{1/ 2}~\cdot \cancel n\cdot (x-2)^{n-1}}{ \cancel n\cdot (n-1)!} \\=\sum_{n=1}^{\infty} \frac{(-1)^{n} ~{n}^{1/ 2} (x-2)^{n-1}}{ (n-1)!} }$$

OpenStudy (perl):

because n! = n (n-1)(n-2)...(3)(2)(1) = n (n-1)!

OpenStudy (anonymous):

Ok I understand that part...

OpenStudy (anonymous):

But I don't understand thomas' explanation...

OpenStudy (perl):

thomas expanded the series itself then take the derivative 7 times, terms will vanish

OpenStudy (anonymous):

Oh I see

OpenStudy (anonymous):

I understand now, thank you :)

OpenStudy (thomas5267):

Or put it this way. \[ \begin{align*} f(x)&=\sum_{n=0}^\infty \frac{(-1)^{n} \sqrt{n}}{ n!}(x-2)^{n}\\ &=\sum_{n=0}^\infty c_n(x-2)^{n}\qquad c_n=\frac{(-1)^n\sqrt{n}}{n!}\\ &=c_1(x-2)+c_2(x-2)^2+c_3(x-3)^3+\cdots\qquad c_0=0\\ f^{(7)}(x)&=7!c_7+(x-2)sth\\ f^{(7)}(2)&=\frac{7!(-1)^7\sqrt{7}}{7!}=-\sqrt{7} \end{align*} \]

OpenStudy (perl):

one thing we might want to do is change the lower limit, because its a bit odd to start at n = 7

OpenStudy (perl):

if we do it the other way we get an indeterminate expression 0^0 $$ \Large { f(x) = \sum_{n=0}^{\infty} \frac{(-1)^{n} ~{n}^{1/ 2}~ (x-2)^{n}}{ n!} \\~\\f^{(7)}(x) = \sum_{n=0}^{\infty} \frac{(-1)^{n+1} ~{(n+7)}^{1/ 2}(x-2)^{n}}{n!} \\~\\f^{(7)}(2) = \sum_{n=0}^{\infty} \frac{(-1)^{n+1} ~{(n+7)}^{1/ 2}(x-2)^{n}}{n!} }$$

OpenStudy (perl):

i'm a bit confused as to which expression it should be, can you take a screenshot of the original problem

OpenStudy (perl):

i guess we can allow that 0^0 = 1 , in this series. for instance we allow it in other cases $$ \Large {e^x = \sum_{n=0}^{\infty } \frac{x^n}{n!} \\~\\ e^0 = \frac{0^0}{0!} + \frac{0^1}{1!} + \frac{0^2}{2!} + ... }$$

OpenStudy (perl):

ok then your work and my work makes sense

OpenStudy (thomas5267):

Can we justify \(0^0\)=1 by saying \(\lim_{x\to 0}x^0=1\)?

OpenStudy (perl):

\[ \Large { f(x) = \sum_{n=0}^{\infty} \frac{(-1)^{n} ~{n}^{1/ 2}~ (x-2)^{n}}{ n!} \\~\\f^{(7)}(x) = \sum_{n=0}^{\infty} \frac{(-1)^{n+1} ~{(n+7)}^{1/ 2}(x-2)^{n}}{n!} \\~\\f^{(7)}(2) = \sum_{n=0}^{\infty} \frac{(-1)^{n+1} ~{(n+7)}^{1/ 2}(x-2)^{n}}{n!} \\~\\ = \frac{(-1)^1\cdot 7^{1/2} \cdot 0^0}{0!} +0 + 0 + ... }\]

OpenStudy (perl):

yes that makes sense, we can plug in the discontinuity

OpenStudy (perl):

but strictly speaking it is indeterminate , i believe

OpenStudy (perl):

what if you did instead lim 0^x , as x->0 , then you would have 0

OpenStudy (perl):

$$ \Large{ \lim_{x\to 0} ~0^x = 0 \\~\\ \lim_{x\to 0}~ x^0 = 1 } $$

OpenStudy (thomas5267):

I think \(0^0\) is not defined precisely because of this reason.

OpenStudy (perl):

but in the series we see that 0^0 has to be equal to 1, as we see in the series e^0

OpenStudy (perl):

\[ \large {e^x = \sum_{n=0}^{\infty } \frac{x^n}{n!} \\~\\ e^0 = \frac{0^0}{0!} + \frac{0^1}{1!} + \frac{0^2}{2!} + ... =\frac{0^0}{1 } + 0 + 0 + ... = \frac{0^0}{1 } \\~\\e^0 = \frac{0^0}{1 } \\~\\1= \frac{0^0}{1 } \\~\\1\cdot 1 = 0^0 }\]

OpenStudy (perl):

we might as well add another limit \[ \Large{ \lim_{x\to 0} ~0^x = 0 \\~\\ \lim_{x\to 0}~ x^0 = 1 \\~\\ \lim_{x\to 0^{+}}~ x^x = 1 }\]

OpenStudy (perl):

interestingly enough my software computes 0^0 as 1 http://prntscr.com/76bw11 but my calculator gives me error

OpenStudy (thomas5267):

Can we just say \(\lim_{(x,y)\to(0,0)}x^y\) does not exist?

OpenStudy (perl):

right, that combines the two cases above

OpenStudy (perl):

and when x = y, the limit is 1

OpenStudy (perl):

wolfram https://www.wolframalpha.com/input/?i=0%5E0

OpenStudy (perl):

err, actually three cases you can use to show limit does not exist, but two different paths are sufficient,